# Trustworthy Transfer Learning: A Survey

### v1

@@ 4.1.2 Federated Transfer
In contrast to the unidirectional hypothesis transfer discussed in the previous subsection, federated transfer emphasizes bidirectional knowledge sharing that allows source and target domains to communicate and exchange information while maintaining privacy protection (?). This is largely inspired by recent personalized federated learning frameworks (?, ?, ?) which allow private clients to collaborate in training personalized models under the coordination of a central server. As illustrated in (?, ?), during each communication round, private clients upload their model updates to the central server, which then securely aggregates these updates and broadcasts the updated model back to each client. In this process, each client exclusively owns their data, which will not be shared with the central server or with other clients. From the perspective of knowledge transferability, the intuition behind personalized federated learning is to transfer knowledge across private clients in a privacy-preserving manner (?, ?, ?). In other words, each private (target) client participates in federated collaboration to receive knowledge from other (source) clients, and its uploaded parameters can be used as indicators to select only the most relevant source knowledge (e.g., related clients with similar parameters within a coalition (?, ?)) under distribution shifts across clients. Formally, given a set of private clients, each with access to a private training set, the problem of federated transfer can be defined as follows.
Definition 8 (Federated Transfer)
@@
$A^{\mathrm{fl}}:(\mathcal{X}\times\mathcal{Y})^{n_{k}}\times\left(\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k-1}\times\mathcal{F}_{k+1}\times\cdots\times\mathcal{F}_{K}\right)\to\mathcal{F}_{k}$ (17)
@@
Note that $\mathcal{F}_{1}=\cdots=\mathcal{F}_{K}$ implies that all clients will share the same hypothesis space. It can be seen from Figure 9 that there are two different scenarios: one where the generalization performance of all clients is important, and another where only the generalization performance of a target client matters.
@@
Figure 9: Illustration of personalized federated learning in two scenarios: (1) Generalization performance is evaluated across all clients, where each client (e.g., client $k$ ($k=1,2,\cdots,$)) is considered as a target client and others as source clients for knowledge transfer. (2) Generalization performance is improved only on a specific target client (i.e., only client $k$).
@@
To address the first scenario, various personalized federated learning frameworks have been proposed, including parameter decoupling (?, ?, ?), model interpolation (?, ?, ?), clustering (?, ?), multi-task learning (?), meta-learning (?), knowledge distillation (?, ?), Bayesian learning (?, ?), etc. Despite the impressive performance of these personalized federated learning frameworks across various applications, it is shown (?, ?) that some clients might suffer from negative transfer in the context of personalized federated learning. It implies that their performance can be worse compared to when they train a model solely on their local data without communicating information with other clients. To mitigate negative transfer issues, INCFL (?) is proposed to maximize the incentivized client participation by dynamically adjusting the aggregation weight assigned to each client. FedCollab (?) optimizes the collaboration structure by clustering clients into non-overlapping coalitions based on their distribution distances and data quantities. Similarly, FEDORA (?) adaptively aggregates relevant source knowledge by considering distribution similarities among clients and regularizes local models when the received knowledge has a positive impact on the generalization performance. DisentAFL (?) uses a two-stage knowledge disentanglement and gating mechanism to enhance positive transfer under complex client heterogeneity, e.g., modality heterogeneity, task heterogeneity, and domain heterogeneity among clients.
@@
To address the second scenario, federated domain adaptation (?, ?, ?) has been studied to transfer knowledge from multiple source clients with sufficient labeled samples to a target client with limited or no labeled samples. Unlike standard personalized federated learning, it focuses only on the generalization performance of the target clients. Specifically, inspired by domain adaptation theory (?), ? (?) derive a weighted error bound for federated domain adaptation. Based on this, the FADA algorithm is proposed to disentangle domain-invariant and domain-specific features for each client and then align the domain-invariant features between source and target clients. Similarly, ? (?) leverage knowledge distillation and BatchNorm Maximum Mean Discrepancy (MMD) to address the distribution gaps between source and target clients. More recently, ? (?) theoretically analyze the connections between the generalization performance and aggregation rules of federated domain adaptation. This finding also results in an auto-weighting scheme for optimal combinations of the source and target gradients. In addition to federated domain adaptation, federated domain generalization aims to train models using source clients and then apply these models to previously unseen target clients (?, ?). Notably, ? (?) propose a federated domain generalization benchmark, highlighting the necessity of evaluation scenarios that involve a large number of private clients, high client heterogeneity, and more realistic data sets.
4.2 Adversarial Robustness
@@
It has been observed (?, ?) that modern machine learning models can be easily fooled by adversarial examples that are perceptibly indistinguishable with respect to clean inputs. This survey focuses on exploring the adversarial robustness of knowledge transfer models under assumptions where distribution shifts occur across domains.
4.2.1 Attacks
@@
Recent efforts have been devoted to understanding the adversarial vulnerability of deep transfer learning techniques (?, ?, ?). In the context of transfer learning, evasion attacks aim to generate adversarial examples to fool the learned transfer learning models on the target domain. Initially, by minimizing the feature representation dissimilarity between adversarial and clean target samples from different classes using only the pre-trained model, ? (?) demonstrate the vulnerability of fine-tuned models in the transfer learning framework. Based on the observations that the neurons of the activation vector within the pre-trained model correlate with target classes, ? (?) design a simple brute-force attacking mechanism. This approach crafts input data to trigger those neurons individually, thereby exploring which one is highly associated with each target class.
@@
Figure 10: Illustration of poisoning attacks on transfer learning. (a) By injecting adversarial noise into the source data, the adversary can control the prediction behavior of transfer learning models on the target domain. One intuitive explanation (?, ?) is that (b) the source and target distributions can be correctly aligned, but (c) they become misaligned after applying the attack.
@@
In contrast, as shown in Figure 10, poisoning attacks allow crafting source samples to control the prediction behavior of transfer learning models on the target domain during model training. Generally, poisoning attacks can occur in two transfer learning scenarios. The first scenario is the joint training of source and target data, assuming that source and target domains have the same labeling space, and the target domain has only unlabeled training samples (i.e., unsupervised domain adaptation (?)). It is motivated by the findings (?) that the feature-based marginal distribution matching can result in negative transfer when the target domain has no label information. Notably, I2Attack (?) and AdaptAttack (?) maximize the label-informed joint distribution discrepancy between the raw source domain and the poisoned source domain with the following constraints. (1) Perceptibly Unnoticeable: All the poisoned input images are natural-looking. (2) Adversarially Indirect: Only source samples are maliciously manipulated. (3) Algorithmically Invisible: Neither source classification error nor marginal domain discrepancy between source and target domains increases. These constraints imply that in the context of transfer learning, an adversary could potentially manipulate the source data to gain control over the prediction function on the target domain. Similarly, ? (?) propose to generate poisoned source samples with clean labels or mislabeled source samples to fool the discrepancy-based adaptation approaches.
@@
The second scenario is the pre-training and fine-tuning framework, where a model is first pre-trained on the source domain and then fine-tuned on the downstream target domain. In this scenario, backdoor attacks are intended to manipulate pre-trained model weights, thus resulting in malicious prediction behavior of fine-tuned models on the target domain (?, ?, ?). The intuition behind backdoor attacks is that when the triggers (e.g., keywords) are activated on target samples, the fine-tuned model will predict pre-defined class labels (?). Specifically, backdoor attacks in pre-trained models satisfy the following conditions (?, ?). (1) Only pre-trained model weights are manipulated, and the infection should be done on the target data through transfer learning. (2) With poisoned pre-trained models, fine-tuned models will behave normally on clean target data, but misclassify any sample with the trigger into a specific class. (3) The designed attacks should be unnoticeable from the viewpoint of the target learner, i.e., the attacker does not alter the fine-tuning process and the training data on the target domain. With these conditions in mind, BadNets (?) directly uses a poisoned data set to adjust the parameters of pre-trained models. RIPPLe (?) poisons the weights of pre-trained models using a bi-level optimization objective over both the poisoning and fine-tuning losses. However, it is noted (?) that the fine-tuned models can mitigate the impact of backdoor attacks during the fine-tuning process on clean target samples. Thus, ? (?) and ? (?) focus on poisoning only the lower layers of pre-trained models. Furthermore, ? (?) reformulate backdoor injection as a lightweight knowledge editing problem and adjust only a subset of model parameters (e.g., key-value pairs) with a minimal amount of poisoned data. Rather than manipulating pre-trained model weights, recent works explore backdoor attacks on large language models (LLMs) in the fine-tuning phase by inserting triggers into instructions (?, ?, ?) or prompts (?, ?, ?). Besides, it is empirically observed (?, ?) that larger LLMs are more susceptible to poisoning attacks than smaller ones. All these backdoor attacks above highlight security and ethical concerns in developing and deploying pre-trained models (?).
4.2.2 Defenses
@@
In the context of transfer learning, the adversarial robustness of the prediction function on the target domain can be improved in various scenarios. (1) Given an adversarially pre-trained source model, the adversarial robustness of source model can be transferred to the target domain (?, ?). (2) Given a standard pre-trained model, the adversarial robustness of the target learner can be enhanced via robust fine-tuning (?, ?). (3) Given an attacked pre-trained model, the defense mechanism can be developed to mitigate the negative impact of source knowledge on the target domain during fine-tuning (?, ?).
@@
Recent works (?, ?, ?) empirically demonstrate the transferability of adversarial robustness across domains, e.g., the robustness of an adversarially pre-trained source model can be transferred to the target domain. Specifically, based on the Learning without Forgetting (LoF) approach (?), ? (?) and ? (?) use the distillation regularization to preserve the robust feature representations of the source model during fine-tuning. The intuition is that the lower layers of an adversarially pre-trained source model can capture robust features from input samples. Similarly, to enhance the transferability of adversarial robustness across domains, ? (?) utilize knowledge distillation to preserve the feature correlations of the robust source model on the target domain. ? (?) propose enforcing feature similarity between natural samples and their corresponding adversarial counterparts during pre-training and regularizing the Lipschitz constant of neural networks during fine-tuning. ? (?) propose a TWINS structure to incorporate the means and variances of batch normalization layers over both pre-training and target data during adversarial fine-tuning. Notably, the studies mentioned above focus on the transferability of empirical adversarial robustness through adversarial training techniques (?, ?), which minimize the adversarial objective against pre-determined strong attacks. In addition to empirical adversarial robustness, ? (?) and ? (?) further investigate certified/provable adversarial robustness (?, ?) in the context of transfer learning, which seeks to maximize the radius around inputs within which the model output remains consistent. Theoretically, ? (?) show that the transferability of adversarial robustness can be guaranteed if the feature extractor of the pre-trained source model is robust and only the newly added linear predictor is updated during fine-tuning. This analysis is consistent with the empirical observations in (?, ?) that feature extractors from an adversarially pre-trained source model contribute to robustness transfer across domains, where only the last layer is re-trained on the target data.
@@
The second line of research is robust fine-tuning (?, ?, ?), where a standard pre-trained model is fine-tuned using adversarial training (?, ?). Specifically, RIFT (?) maximizes the mutual information between the feature extracted by the adversarially fine-tuned model and the class label plus the feature extracted by the pre-trained model. AutoLoRa (?) disentangles robust fine-tuning via a low-rank branch to mitigate gradient conflicts between adversarial and natural objectives. It optimizes adversarial objective w.r.t. the standard feature extractor and standard objective w.r.t. the auxiliary LoRA branch. More recently, ? (?) examine the adversarial robustness of hypothesis transfer learning (?), which involves transferring knowledge from source domains to a target domain using a set of pre-trained auxiliary hypotheses. They derive generalization error bounds for adversarial robustness in the target domain based on two specific algorithms: adversarial regularized empirical risk minimization and proximal stochastic adversarial training.
@@
The previous two scenarios assume the availability of a clean pre-trained source model for transfer learning. Their goal is to improve the adversarial robustness of fine-tuned models against adversarial perturbations in the target samples during inference. As discussed in Subsection 4.2.1, a more challenging yet realistic scenario occurs when a poisoned source model (?, ?) is deployed for transfer learning. In this scenario, defense mechanisms should handle the negative impact of poisoned source knowledge during fine-tuning. To this end, ? (?) design a defense mechanism to counter the attack proposed by ? (?). The key idea is to reduce the similarity between the pre-trained and the fine-tuned models via noisy feature distillation. More recently, in the context of backdoored large language models (LLMs) (?, ?), ? (?) propose to detect poisoned target samples associated with triggers during inference by leveraging the different masking-sensitivity of poisoned and clean samples. The intuition is that poisoned samples are more sensitive to random masking than clean samples, as fine-tuned LLMs might exhibit significant changes in predictions when the trigger and normal content are masked within poisoned samples. Similarly, ? (?) and ? (?) detect poisoned samples using the perplexity changes of samples under word deletion or different robustness properties of clean and poisoned samples against triggers, respectively.
4.2.3 Transferability vs. Robustness
@@
In addition to highlighting the adversarial vulnerability and robustness of transfer learning frameworks, recent studies have also explored the connection between knowledge transferability and adversarial robustness (?, ?). To be specific, ? (?) empirically demonstrate that adversarially robust models can transfer better (i.e., higher transfer accuracy in the target domain) than their standard-trained counterparts. That is, though robustness may be at odds with accuracy within the same domain (?), the adversarial robustness achieved in a source domain can improve the transfer accuracy in a related target domain. ? (?) further explain that in image classification tasks, adversarial training in the source domain biases the learned representations towards retaining shapes, thereby improving transferability in the target domain. Theoretically, ? (?) provide an information-theoretic justification for adversarial training, implying the trade-off between accuracy on the source domain and transferability on a related target domain. More rigorously, ? (?) demonstrate that, for a learning function based on a two-layer linear neural network, adversarially robust representation learning over multiple source domains leads to much tighter transfer error bounds on the target domain than standard representation learning. Alternatively, ? (?) argue that adversarial training regularizes the function class of feature representation learning, thus enhancing knowledge transferability across domains.
4.3 Fairness

### v2

@@ 4.1.2. Federated Transfer
In contrast to the unidirectional hypothesis transfer discussed in the previous subsection, federated transfer emphasizes bidirectional knowledge sharing that allows source and target domains to communicate and exchange information while maintaining privacy protection (FedAvg). This is largely inspired by recent personalized federated learning frameworks (KairouzMABBBBCC21; liu2020secure) which allow private clients to collaborate in training personalized models under the coordination of a central server. As illustrated in (FedAvg; KairouzMABBBBCC21), during each communication round, private clients upload their model updates to the central server, which then securely aggregates these updates and broadcasts the updated model back to each client. In this process, each client exclusively owns their data, which will not be shared with the central server or with other clients. From the perspective of knowledge transferability, the intuition behind personalized federated learning is to transfer knowledge across private clients in a privacy-preserving manner (FedHealth; FEDORA; liu2020secure). In other words, each private (target) client participates in federated collaboration to receive knowledge from other (source) clients, and its uploaded parameters can be used as indicators to select only the most relevant source knowledge (e.g., related clients with similar parameters within a coalition (FedCollab; donahue2021model)) under distribution shifts across clients. Formally, given a set of private clients, each with access to a private training set, the problem of federated transfer can be defined as follows.
Definition 4.2 (Federated Transfer).
@@
(17) $A^{\mathrm{fl}}:(\mathcal{X}\times\mathcal{Y})^{n_{k}}\times\left(\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k-1}\times\mathcal{F}_{k+1}\times\cdots\times\mathcal{F}_{K}\right)\to\mathcal{F}_{k}$
@@
Note that $\mathcal{F}_{1}=\cdots=\mathcal{F}_{K}$ implies that all clients will share the same hypothesis space. It can be seen from Fig. 9 that there are two different scenarios: one where the generalization performance of all clients is important, and another where only the generalization performance of a target client matters.
@@
Figure 9. Illustration of personalized federated learning in two scenarios: (1) Generalization performance is evaluated across all clients, where each client (e.g., client $k$ ($k=1,2,\cdots,$)) is considered as a target client and others as source clients for knowledge transfer. (2) Generalization performance is improved only on a specific target client (i.e., only client $k$).
@@
To address the first scenario, various personalized federated learning frameworks have been proposed, including parameter decoupling (FedRep), model interpolation (Ditto; pFedMe), clustering (CFL; GhoshCYR20), multi-task learning (SmithCST17), meta-learning (Per-FedAvg), knowledge distillation (zhang2021parameterized; zhu2021data), Bayesian learning (achituve2021personalized; zhang2022personalized), etc. Despite the impressive performance of these personalized federated learning frameworks across various applications, it is shown (FEDORA; FedCollab) that some clients might suffer from negative transfer in the context of personalized federated learning. It implies that their performance can be worse compared to when they train a model solely on their local data without communicating information with other clients. To mitigate negative transfer issues, INCFL (cho2022to) is proposed to maximize the incentivized client participation by dynamically adjusting the aggregation weight assigned to each client. FedCollab (FedCollab) optimizes the collaboration structure by clustering clients into non-overlapping coalitions based on their distribution distances and data quantities. Similarly, FEDORA (FEDORA) adaptively aggregates relevant source knowledge by considering distribution similarities among clients and regularizes local models when the received knowledge has a positive impact on the generalization performance. DisentAFL (chen2024disentanglement) uses a two-stage knowledge disentanglement and gating mechanism to enhance positive transfer under complex client heterogeneity, e.g., modality heterogeneity, task heterogeneity, and domain heterogeneity among clients.
@@
To address the second scenario, federated domain adaptation (PengHZS20; Fan0DH23; jiang2024principled) has been studied to transfer knowledge from multiple source clients with sufficient labeled samples to a target client with limited or no labeled samples. Unlike standard personalized federated learning, it focuses only on the generalization performance of the target clients. Specifically, inspired by domain adaptation theory (A_distance), (PengHZS20) derive a weighted error bound for federated domain adaptation. Based on this, the FADA algorithm is proposed to disentangle domain-invariant and domain-specific features for each client and then align the domain-invariant features between source and target clients. Similarly, (feng2021kd3a) leverage knowledge distillation and BatchNorm Maximum Mean Discrepancy (MMD) to address the distribution gaps between source and target clients. More recently, (jiang2024principled) theoretically analyze the connections between the generalization performance and aggregation rules of federated domain adaptation. This finding also results in an auto-weighting scheme for optimal combinations of the source and target gradients. In addition to federated domain adaptation, federated domain generalization aims to train models using source clients and then apply these models to previously unseen target clients (nguyen2022fedsr; yuan2022what). Notably, (bai2024benchmarking) propose a federated domain generalization benchmark, highlighting the necessity of evaluation scenarios that involve a large number of private clients, high client heterogeneity, and more realistic data sets.
4.2. Adversarial Robustness
@@
It has been observed (SzegedyZSBEGF13; GoodfellowSS14) that modern machine learning models can be easily fooled by adversarial examples that are perceptibly indistinguishable with respect to clean inputs. This survey focuses on exploring the adversarial robustness of knowledge transfer models under assumptions where distribution shifts occur across domains.
4.2.1. Attacks
@@
Recent efforts have been devoted to understanding the adversarial vulnerability of deep transfer learning techniques (WangYVZZ18; ZhangSLB020; Rezaei2020A). In the context of transfer learning, evasion attacks aim to generate adversarial examples to fool the learned transfer learning models on the target domain. Initially, by minimizing the feature representation dissimilarity between adversarial and clean target samples from different classes using only the pre-trained model, (WangYVZZ18) demonstrate the vulnerability of fine-tuned models in the transfer learning framework. Based on the observations that the neurons of the activation vector within the pre-trained model correlate with target classes, (Rezaei2020A) design a simple brute-force attacking mechanism. This approach crafts input data to trigger those neurons individually, thereby exploring which one is highly associated with each target class.
(a) Poisoning attacks on transfer learning
(b) Before attack
(c) After attack
@@
Figure 10. Illustration of poisoning attacks on transfer learning. (a) By injecting adversarial noise into the source data, the adversary can control the prediction behavior of transfer learning models on the target domain. One intuitive explanation (I2Attack; MehraKCH21) is that (b) initially the source and target distributions can be correctly aligned, but (c) they become misaligned after applying the attack.
@@
In contrast, as shown in Fig. 10, poisoning attacks allow crafting source samples to control the prediction behavior of transfer learning models on the target domain during model training. Generally, poisoning attacks can occur in two transfer learning scenarios. The first scenario is the joint training of source and target data, assuming that source and target domains have the same labeling space, and the target domain has only unlabeled training samples (i.e., unsupervised domain adaptation (TL_survey_2010)). It is motivated by the findings (0002CZG19) that the feature-based marginal distribution matching can result in negative transfer when the target domain has no label information. Notably, I2Attack (I2Attack) and AdaptAttack (AdaptAttack) maximize the label-informed joint distribution discrepancy between the raw source domain and the poisoned source domain with the following constraints. (1) Perceptibly Unnoticeable: All the poisoned input images are natural-looking. (2) Adversarially Indirect: Only source samples are maliciously manipulated. (3) Algorithmically Invisible: Neither source classification error nor marginal domain discrepancy between source and target domains increases. These constraints imply that in the context of transfer learning, an adversary could potentially manipulate the source data to gain control over the prediction function on the target domain. Similarly, (MehraKCH21) propose to generate poisoned source samples with clean labels or mislabeled source samples to fool the discrepancy-based adaptation approaches.
@@
The second scenario is the pre-training and fine-tuning framework, where a model is first pre-trained on the source domain and then fine-tuned on the downstream target domain. In this scenario, backdoor attacks are intended to manipulate pre-trained model weights, thus resulting in malicious prediction behavior of fine-tuned models on the target domain (JiZJLW18; ShenJ0LCSFYW21). The intuition behind backdoor attacks is that when the triggers (e.g., keywords) are activated on target samples, the fine-tuned model will predict pre-defined class labels (BadNets). Specifically, backdoor attacks in pre-trained models satisfy the following conditions (RIPPLe; YaoLZZ19). (1) Only pre-trained model weights are manipulated, and the infection should be done on the target data through transfer learning. (2) With poisoned pre-trained models, fine-tuned models will behave normally on clean target data, but misclassify any sample with the trigger into a specific class. (3) The designed attacks should be unnoticeable from the viewpoint of the target learner, i.e., the attacker does not alter the fine-tuning process and the training data on the target domain. With these conditions in mind, BadNets (BadNets) directly uses a poisoned data set to adjust the parameters of pre-trained models. RIPPLe (RIPPLe) poisons the weights of pre-trained models using a bi-level optimization objective over both the poisoning and fine-tuning losses. However, it is noticed (RIPPLe) that the fine-tuned models can mitigate the impact of backdoor attacks during the fine-tuning process on clean target samples. Thus, (YaoLZZ19; LiSLZMQ21) focus on poisoning only the lower layers of pre-trained models. Furthermore, (BadEdit) reformulate backdoor injection as a lightweight knowledge editing problem and adjusts only a subset of model parameters (e.g., key-value pairs) with a minimal amount of poisoned data. Rather than manipulating pre-trained model weights, recent works explore backdoor attacks on large language models (LLMs) in the fine-tuning phase by inserting triggers into instructions (xu2024instructions; yan2024backdooring; ShuW0GXG23) or prompts (jiang2023forcing; ZhaoWLZF23; BadPrompt). Besides, it is empirically observed (WanWSK23; bowen2024scaling) that larger LLMs are more susceptible to poisoning attacks than smaller ones. All these backdoor attacks above highlight security and ethical concerns in developing and deploying pre-trained models (hubinger2024sleeper).
4.2.2. Defenses
@@
In the context of transfer learning, the adversarial robustness of the prediction function on the target domain can be improved in various scenarios. (1) Given an adversarially pre-trained source model, the adversarial robustness of the source model can be transferred to the target domain (HendrycksLM19; ShafahiSZGSJG20). (2) Given a standard pre-trained model, the adversarial robustness of the target learner can be enhanced via robust fine-tuning (AutoLoRa). (3) Given an attacked pre-trained model, the defense mechanism can be developed to mitigate the negative impact of source knowledge on the target domain during fine-tuning (ChinZM21; MDP).
@@
Recent works (davchev2019empirical; HendrycksLM19; Chen0C0AW20) empirically demonstrate the transferability of adversarial robustness across domains, e.g., the robustness of an adversarially pre-trained source model can be transferred to the target domain. Specifically, based on the Learning without Forgetting (LoF) approach (li2017learning), (ShafahiSZGSJG20; VaishnaviER22) use the distillation regularization to preserve the robust feature representations of the source model during fine-tuning. The intuition is that the lower layers of an adversarially pre-trained source model can capture robust features from input samples. Similarly, to enhance the transferability of adversarial robustness across domains, (RFA) utilize knowledge distillation to preserve the feature correlations of the robust source model on the target domain. (CARTL) propose enforcing feature similarity between natural samples and their corresponding adversarial counterparts during pre-training and regularizing the Lipschitz constant of neural networks during fine-tuning. (TWINS) propose a TWINS structure to incorporate the means and variances of batch normalization layers over both pre-training and target data during adversarial fine-tuning. Notably, the studies mentioned above focus on the transferability of empirical adversarial robustness through adversarial training techniques (MadryMSTV18; GoodfellowSS14), which minimize the adversarial objective against pre-determined strong attacks. In addition to empirical adversarial robustness, (alhamoud2023generalizability; vaishnavi2024a) further investigate certified/provable adversarial robustness (CohenRK19; JeongS20) in the context of transfer learning, which seeks to maximize the radius around inputs within which the model output remains consistent. Theoretically, (NernRGS23) show that the transferability of adversarial robustness can be guaranteed if the feature extractor of the pre-trained source model is robust and only the newly added linear predictor is updated during fine-tuning. This analysis is consistent with the empirical observations (ShafahiSZGSJG20; hua2024initialization) that feature extractors from an adversarially pre-trained source model contribute to robustness transfer across domains, where only the last layer is re-trained on the target data.
@@
The second line of research is robust fine-tuning (DongLLYZ21; AutoLoRa), where a standard pre-trained model is fine-tuned using adversarial training (MadryMSTV18; GoodfellowSS14). Specifically, RIFT (DongLLYZ21) maximizes the mutual information between the feature extracted by the adversarially fine-tuned model and the class label plus the feature extracted by the pre-trained model. AutoLoRa (AutoLoRa) disentangles robust fine-tuning via a low-rank branch to mitigate gradient conflicts between adversarial and natural objectives. It optimizes adversarial objective w.r.t. the standard feature extractor and standard objective w.r.t. the auxiliary LoRA branch. More recently, (wang2024adversarially) examine the adversarial robustness of hypothesis transfer learning (KuzborskijO13), which involves transferring knowledge from source domains to a target domain using a set of pre-trained auxiliary hypotheses. They derive generalization error bounds for adversarial robustness in the target domain based on two specific algorithms: adversarial regularized empirical risk minimization and proximal stochastic adversarial training.
@@
The previous two scenarios assume the availability of a clean pre-trained source model for transfer learning. Their goal is to improve the adversarial robustness of fine-tuned models against adversarial perturbations in the target samples during inference. As discussed in Subsection 4.2.1, a more challenging yet realistic scenario occurs when a poisoned source model (Rezaei2020A; BadPrompt) is deployed for transfer learning. In this scenario, defense mechanisms should handle the negative impact of poisoned source knowledge during fine-tuning. To this end, (ChinZM21) design a defense mechanism to counter the attack proposed by (Rezaei2020A). The key idea is to reduce the similarity between the pre-trained and the fine-tuned models via noisy feature distillation. More recently, in the context of backdoored large language models (LLMs) (BadPrompt; ZhaoWLZF23), (MDP) propose to detect poisoned target samples associated with triggers during inference by leveraging the different masking-sensitivity of poisoned and clean samples. The intuition is that poisoned samples are more sensitive to random masking than clean samples, as fine-tuned LLMs might exhibit significant changes in predictions when the trigger and normal content are masked within poisoned samples. Similarly, (QiCLYLS21; YangLLZS21) detect poisoned samples using the perplexity changes of samples under word deletion or different robustness properties of clean and poisoned samples against triggers, respectively.
4.2.3. Transferability vs. Robustness
@@
In addition to highlighting the adversarial vulnerability and robustness of transfer learning frameworks, recent studies have also explored the connection between knowledge transferability and adversarial robustness (SalmanIEKM20; TerziAMS21). To be specific, it is empirically demonstrated (SalmanIEKM20) that adversarially robust models can transfer better (i.e., higher transfer accuracy in the target domain) than their standard-trained counterparts. That is, though robustness may be at odds with accuracy within the same domain (TsiprasSETM19), the adversarial robustness achieved in a source domain can improve the transfer accuracy in a related target domain. (UtreraKEKM21) further explain that in image classification tasks, adversarial training in the source domain biases the learned representations towards retaining shapes, thereby improving transferability in the target domain. Theoretically, (TerziAMS21) provide an information-theoretic justification for adversarial training, implying the trade-off between accuracy on the source domain and transferability on a related target domain. More rigorously, (DengZVKZ21) demonstrate that, for a learning function based on a two-layer linear neural network, adversarially robust representation learning over multiple source domains leads to much tighter transfer error bounds on the target domain than standard representation learning. Alternatively, (XuZMSKL22) show that adversarial training regularizes the function class of feature representation learning, thus improving knowledge transferability across domains.
4.3. Fairness